the position x of a particle along straight line is given as x = 3t^2-18t. the max speed in time interval t=0 to t=2 sec is?

Grade:

2 Answers

Nicho priyatham
8 years ago
given x=3t2-18t
so since v=dx/dt
we get v=6t-18
so in intrewal 0
so vmax=6(2)-18=-6 (neglect -sign coz it indicates direction)
so v max is 6
PLZZ APPROVE if useful and in case of any doubts put it in ans box
28B Mahima Nathan
2 years ago
The answer is 18m/s. This is because velocity will be maximum at 0s. 
v=(0*6)-18= -18 and since – just indicates direction 18m/s is the max velocity.